OpenStudy (anonymous):

given the function F(x)= -5|x+1|+3 for what values of x is F(x)=-12 my choices are A) x=-2,x=-4 B) x=-2,x=4 C)x=2,x=-4 D)x=2,x=4

imqwerty (imqwerty):

okay so we have this- \(\large F(x)=-5|x+1|+3\) we gotta find x for which \(F(x)\) is -12 so \(\large -12=-5|x+1|+3\) \(\large -12-3=-5|x+1|+3-3\) \(\large -15=-5|x+1|\) \(\huge\frac{-15}{-5}=\frac{-5|x+1|}{-5}\) \(\large 3=|x+1|\) so either \(x+1=3 \) or \(x+1=-3\) solving these two we get the value of \(x\) to be 2 and -4

OpenStudy (anonymous):

awesome thanks

imqwerty (imqwerty):

np :)